OpenStudy (anonymous):

Derive the equation of the parabola with a focus at (−7, 5) and a directrix of y = −11 Answer choices: a) f(x) = 1/32 (x+7)^2-3 b) f(x) = -1/32 (x+7)^2-3 c) f(x) = -1/32 (x-7)^2 -3 d) f(x) = 1/32 (x-7)^2 -3

OpenStudy (michele_laino):

Please, if (x,y) is a point of your parabola, what is the distance between (x,y) and the focus (-7,5)?

OpenStudy (anonymous):

(-7, -6 ) ?

OpenStudy (anonymous):

oops wait i mean ( -7, -3 )?

OpenStudy (michele_laino):

you ahve to apply this formula, distance d is: \[d=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y _{1})^{2}}\] where (x_1,y_1) is the initial point and (x_2,y_2) is the focus

OpenStudy (michele_laino):

namely, (x_1,y_1)=(x,y) and (x_2,y_2)=(-7,5). Please substitute them into the formula for d

OpenStudy (michele_laino):

for example, we have: \[x _{2}-x _{1}=x-(-7)=x+7\] and: \[y _{2}-y _{1}=y-5\] so, please substitute them

OpenStudy (anonymous):

ok so (x+7)^2 = x^2 + 14x + 49 and (y-5)^2 = y^2 -10y +25

OpenStudy (michele_laino):

ok!, so: \[d=\sqrt{x ^{2}+49+14x+y ^{2}+15-10y}\]

OpenStudy (anonymous):

is the answer x^2 + y^2 + 14x - 10y _ 64?

OpenStudy (michele_laino):

Now, what is the distance between the point (x,y) and the directrix? please note this drawing: